OpenStudy (anonymous):

please help and show steps: x^2 + x -56

OpenStudy (anonymous):

\[x ^{2}+x-56\]

OpenStudy (anonymous):

assume this = 0

OpenStudy (ybarrap):

What two numbers multiply to give you -56 but add to give you 1?

OpenStudy (ybarrap):

(x+a)(x+b) ab=-56 a+b=1 a=-7 b=8

OpenStudy (ybarrap):

So, $$ \large x^2+x-56=(x+8)(x-7) $$